OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5/9 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 20 degrees Celsius is

OpenStudy (anonymous):

First, multiply both sides by 9/5 (to make the 5/9 on the right equal 1) 9/5 C = F-32 Then add 32 two both sides, you get 9/5C - 32 = F Plug in 20 for C and you are set